# A. I. Bolotnikov, A. A. Irmatov

### v1

@@ Introduction
In [1] the matching arrangement of hyperplanes $MA(G)$ for a graph $G(V,E),\\ |E|=n$ without loops and parallel edges was introduced as the arrangement that contains a hyperplane with the equation $x_{i_{1}}-x_{i_{2}}+...+(-1)^{k+1}x_{i_{k}}=0$ for every sequence of edges $(e_{i_{1}},...,e_{i_{k}})$ that form either a simple path of arbitrary length or a simple cycle of even length in $G$. The matching arrangement has the following property: if you arbitrarily select two vectors from the same region of the arrangement, use each vector in turn as a weight vector on the edges of graph $G$, and find the maximum-weight matching for each vector, then for both vectors, the maximum matching will be the same [1]. In addition, there is a one-to-one correspondence between the regions of the matching arrangement and the LP-orientations of the matching polytope, a special polytope used in solving the maximum matching problem using linear programming methods [2].
In [3], Thomas Zaslavsky showed that the number of regions in a hyperplane arrangement depends on the partially ordered set of intersections of the hyperplanes in the arrangement. Zaslavsky expressed this number through the characteristic polynomial of the hyperplane arrangement. Thus, if we calculate the characteristic polynomial of a hyperplane arrangement, we can find its number of regions.
@@
Let A be any subspace arrangement in $R^{n}$ defined over the integers and q be a large enough prime number. Then $\chi_{A}(q)=|F_{q}^{n}\setminus\cup A|$ Equivalently, identifying $F_{q}^{n}$ with $\{0,1,...,q-1\}^{n}$, $\chi_{A}(q)$ is the number of points with integer coordinates in the cube $\{0,1,...,q-1\}^{n}$ which do not satisfy mod q the defining equations of any of the subspaces in A.
The second section of this paper presents results on calculating the characteristic polynomials of the matching arrangement for some families of graphs using the finite field method.The definitions of a hyperplane arrangement, an intersection lattice and a characteristic polynomial that were used in the first part can be found in [6], which also contains the proof of the deletion-restriction lemma and the description of the finite field method that uses powers of prime.
@@
The forth section of the paper provides a theoretical description of the Bolotnikov-Irmatov cryptosysyem that uses the NP-completeness of the improper weight function problem.
The results from the second and the third sections are obtained by A. I. Bolotnikov. The cryptosystem in the forth section is developed by A. I. Bolotnikov and A. A. Irmatov.
The characteristic polynomial of the matching arrangement for some classes of graphs.
@@
Let $G$ be a graph that consists of one cycle of length $n$, and let $\chi_{MA(G)}(x)$ be the characteristic polynomial of the matching arrangement $MA(G)$. Then
@@
Let graph G consist of one cycle of length n. We arbitrarily choose vertex A in graph G and number the edges clockwise, counting the edges from vertex A.
@@
Let us first consider the case when n is even. Consider the linear operator $W:R^{n}\rightarrow R^{n}$ with the following matrix. The main diagonal of the matrix, as well as the next diagonal to the right, alternate between 1 and -1 (starting with 1), and there are 0s in all other places.
@@
The vector $t=(1,-1,...,(-1)^{k+1},0,...,0)$ is the normal vector to the hyperplane from the arrangement $MA(G)$, which corresponds to a path of length k leaving A and going clockwise. The operator W maps this vector to vector $d_{k}=(0,...,0,1,0,...,0)$, in which 1 is at the $k$-th position. For an arbitrary vertex B of graph G, the vector $t_{B}=(1,-1,...,(-1)^{r+1},0,...,0)$, where r is the length of the path from A to B clockwise, is normal to the hyperplane corresponding to the path from A to B in a clockwise direction. $t_{A}=(1,-1,...,1,-1)$ will denote the normal to the hyperplane corresponding to the entire cycle. Consider an arbitrary path in graph G that leads clockwise from vertex B to vertex C. If $B=A$, then the vector $t_{C}$ will be normal to the hyperplane corresponding to this path. If $C=A$ or the path does not pass through A, then $t_{C}-t_{B}$ will be normal to the corresponding hyperplane. If the path passes through A but does not start or end there, then $t_{A}-t_{B}+t_{C}$ will be normal to the corresponding hyperplane. Operator W maps $t_{B}$ to $d_{r}$, $t_{C}-t_{B}$ to $d_{k_{2}}-d_{k_{1}}$, where $k_{2}>k_{1}$, and $t_{A}-t_{B}+t_{C}$ to $d_{n}-d_{k_{1}}+d_{k_{2}}$, where $k_{1}>k_{2}$.
@@
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}+x_{i_{1}}=0$
Note that since there is an invertible linear operator W that maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of S , their intersection lattices are isomorphic and their characteristic polynomials are equal. We will look for the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
@@
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}+x_{i_{1}}=0$.
We will find the number $\chi_{S}(p)$ combinatorially. First, we choose a value for $x_{n}$. This can be done in $p-1$ ways. Let’s say we have chosen $x_{n}=k$, $k\neq 0$. Then the remaining $x_{i}$ should be selected so that they do not satisfy any of the following conditions:
@@
Let’s choose a weak ordered partition $\pi=(B_{1},B_{2},....,B_{p-n})$ of the set $\{1,2,3,..,n-1\}$ into $p-n$ blocks (“weak” means that some of the blocks can be empty), that is $\bigcup B_{i}=\{1,2,...,n\}$ and $B_{i}\cap B_{j}=\emptyset,i\neq j$. Arrange all the elements of $F_{p}$ in a row as follows: $0,k,2k,...,(p-1)k$. The first two elements of this row are “locked”: none of the $x_{i}$ can take these two values. Starting with the third element in this row, elements of $B_{1}$ are placed one by one in the descending order onto the elements of the row without any gaps. Once all elements of $B_{1}$ are placed, a gap with the size of one element of the row is made, and elements of $B_{2}$ are placed on the folowing elements of the row in the same descending order. We keep placing elements of $B_{i}$ onto the elements of the row in the descending order and making one-element gaps when moving from $B_{i}$ to $B_{i+1}$, until all elements from $\{1,2,3,...,n-1\}$ are placed. Now for every i, $10\leq i\leq n-$ 1, we choose the element of the row, in which we placed i, as $x_{i}$. Conditions $x_{i}=0$ and $x_{i}=k$ aren’t satisfied for any i, because the first two elements of the row were left “locked”. Condition $x_{i_{2}}-x_{i_{1}}=0$ is not satisfied for any $i_{1}$ and $i_{2}$, because we put no more then one element of $\{1,2,3,4,..,n-1\}$ onto each element of the row. Condition $0\leq i_{1}\leq i_{2}\leq_{n}-1:x_{i_{2}}-x_{i_{1}}=k$ is not satisfied for any $i_{1}\leq i_{2}$, because for any $B_{i}$ the elements of $B_{i}$ were placed in the descending order, and there was a one-element gap between elements of $B_{i}$ and $B_{i+1}$. Therefore, a vector $(x_{1},x_{2},...,x_{n-1})$ is constructed for a given partition $\pi$, and it does not satisfy any of the four conditions. Conversely, given a vector $(x_{1},x_{2},...,x_{n-1})$ that does not satisfy any of the four conditions, a placement of $\{1,2,3,...,n-1\}$ onto the row $0,k,2k,...,(p-1)k$ can be recovered, and in turn a partition $\pi$ can be recovered from this placement. This means that the number of vectors that do not satisfy any of the four conditions is equal to the number of weak ordered partitions of $\{1,2,3,...,n-1\}$ into $p-n$ blocks. There are exactly $(p-n)^{n-1}$ of such partitions. And considering the choice of $x_{n}$ on the first step, the resulting formula for the characteristic polynomial of S is $\chi_{S}(p)=(p-1)(p-n)^{n-1}$. Thus, in case when graph G consists of one cycle of even length, $\chi_{MA(G)}(x)=(x-1)(x-n)^{n-1}$.
Now let n be odd. Let $L$ be the following arrangement of hyperplanes: $L=MA(G)\cup\{H\}$, where H is a hyperplane with the equation $x_{1}-x_{2}+x_{3}...-x_{n-1}+x_{n}=0$.
Consider the linear operator $W:R^{n}\rightarrow R^{n}$ with the following matrix. The main diagonal of the matrix, as well as the next diagonal to the right, alternate between 1 and -1 (starting with 1), and there are 0s in all other places.
The vector $t=(1,-1,...,(-1)^{k+1},0,...,0)$ is the normal vector to the hyperplane from the arrangement $L$, which corresponds to a path of length k leaving A and going clockwise. The operator W maps this vector to vector $d_{k}=(0,...,0,1,0,...,0)$, in which 1 is at the $k$-th position. For an arbitrary vertex B of graph G, the vector $t_{B}=(1,-1,...,(-1)^{r+1},0,...,0)$, where r is the length of the path from A to B clockwise, is normal to the hyperplane corresponding to the path from A to B in a clockwise direction. $t_{A}=(1,-1,...,1,-1,1)$ will denote the normal to $H$. Consider an arbitrary path in graph G that leads clockwise from vertex B to vertex C. If $B=A$, then the vector $t_{C}$ will be normal to the hyperplane corresponding to this path. If $C=A$ or the path does not pass through A, then $t_{C}-t_{B}$ will be normal to the corresponding hyperplane. If the path passes through A but does not start or end there, then $t_{A}-t_{B}-t_{C}$ will be normal to the corresponding hyperplane. Operator W maps $t_{B}$ to $d_{r}$, $t_{C}-t_{B}$ to $d_{k_{2}}-d_{k_{1}}$, where $k_{2}>k_{1}$, and $t_{A}-t_{B}-t_{C}$ to $d_{n}-d_{k_{1}}-d_{k_{2}}$, where $k_{1}>k_{2}$.
@@
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}-x_{i_{1}}=0$
Note that since there is an invertible linear operator W that maps the set of normal vectors to hyperplanes of $L$ to the set of normal vectors to hyperplanes of S , their intersection lattices are isomorphic and their characteristic polynomials are equal. We will look for the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
@@
    $\forall 1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}-x_{i_{1}}=0$.
We will find the number $\chi_{S}(p)$ combinatorially. First, we choose a value for $x_{n}$. This can be done in $p-1$ ways. Let’s say we have chosen $x_{n}=k$, $k\neq 0$. Then the remaining $x_{i}$ should be selected so that they do not satisfy any of the following conditions:
@@
    $\forall 1\leq i_{1}\leq i_{2}\leq_{n}-1:x_{i_{2}}+x_{i_{1}}=k$.
Let us consider two possible cases:
Case 1: $\forall 0\leq i\leq n-1x_{i}\neq k/2$. In this case we can consequitively choose values for all $x_{i}$. There are $p-3$ possible values for $x_{1}$. After a value is chosen for $x_{1}$, there are $k-5$ possible values for $x_{2}$: if value t is chosen for $x_{1}$, then we cannot choose t and $k-t$ for $x_{2}$. After a value is chosen for $x_{2}$,there are $p-7$ possible values for $x_{3}$, and so on. In total in this case there are $(p-3)(p-5)(p-7)...(p-(2n-1))$ possible ways to choose values for $x_{1},x_{2},...,x_{n_{1}}$.
Case 2: for some i $x_{i}=k/2$. In this case we start with choosing in n-1 ways, which of $x_{i}$ is equal to $k/2$, and then for the rest of $x_{i}$ the values are chosen the same way, as in case 1.Therefore, in this case there are $(n-1)(p-3)(p-5)...(p-(2n-3))$ ways to choose values for $x_{1},x_{2},...,x_{n_{1}}$.
@@
Let $\hat{S}$ be an arrangement $\hat{S}=S\setminus{\hat{H}}$, where $\hat{H}$ is an arrangement with the equation $x_{n}=0$. An invertible linear operator W maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of $\hat{S}$ , which means that their intersection lattices are isomorphic and their characteristic polynomials are equal.
The characteristic polynomial of $\hat{S}$ can be found with the use of the deletion-restriction lemma: $\chi_{\hat{S}}(x)=\chi_{S}(x)+\chi_{S_{\hat{H}}}(x)$, where $S_{\hat{H}}$ is an arrangement of hyperplanes in $\hat{H}$ that consists of intersections of $\hat{H}$ with other hyperplanes of S. The following are the equations of hyperplanes of $S_{\hat{H}}$:
@@
    $\forall 1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}+x_{i_{1}}=0$.
We will find $\chi_{S_{\hat{H}}}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations above. Let us choose values for $x_{i}$ consequitively. There are $p-1$ ways to choose a value for $x_{1}$, there are $p-3$ ways to choose a value for $x_{2}$ after choosing a value for $x_{1}$, and so on. Therefore, $\chi_{S_{\hat{H}}}(p)=(p-1)(p-3)(p-5)...(p-(2n-3))$.
After applying the deletion-restriction lemma, the following result is obtained: in case when G consists of one cycle of odd length, $\chi_{MA(G)(x)}=(x-1)(x-3)(x-5)...(x-(2n-3))(x-(n-1))$.
@@
Let graph G consist of two subgraphs that intersect in one vertex - one subgraph is a complete graph $K_{3}$ and the other is the “tail” - a path of an arbitrary length. Then $\chi_{MA(G)}(x)=(x-1)(x-3)(x-4)....(x-(n-2))(x-(n-1))(x-(n-1))(x-n)$.
@@
Let graph G consist of two subgraphs that intersect in one vertex - one subgraph is a complete graph $K_{3}$ and the other is the “tail” - a path of an arbitrary length. Let us number the edges of G in the following order: first we number the edges of $K_{3}$, then we number the edges of the tail, starting with the edge that is adjacent to $K_{3}$.
@@
Let W be a linear operator $W:R^{n}\rightarrow R^{n}$ with the following matrix:
@@
Let $d_{1}=(1,0,...,0)$, $d_{2}=(0,1,0,...,0)$, $d_{3}=(0,1,-1,0,...,0)$,
$d_{i}=(0,0,0,1,-1,...,(-1)^{i},0,...,0),i>3$. Operator W maps $d_{i}$ into a vector $(0,...,0,1,0,...,0)$, where 1 stands on the i-th place.
Let’s consider three possible cases for paths in G.
@@
Case 3: The path goes from $K_{3}$ into the tail. The following are all the possible cases of the intersection of this path and the $K_{3}$ subgraph. If this intersection is $\{e_{1}\}$,then $d_{i}-d_{1}$ is a normal vector to the corresponding hyperplane. If $\{e_{2}\}$ is the intersection, then $d_{i}-d_{2}$ will be the normal vector. If $\{e_{2},e_{3}\}$ is the intersection, then $d_{i}-d_{3}$ will be the normal vector.If $\{e_{3},e_{1}\}$ is the intersection, then $d_{i}-d_{1}+d_{2}-d_{3}$ is the normal vector.
Let S be an arrangement that consists of the following hyperplanes:
@@
    $x_{1}-x_{2}+x_{3}=0$
Note that since there is an invertible linear operator W that maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of S , their intersection lattices are isomorphic and their characteristic polynomials are equal. We will look for the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number p, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},x_{2},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
@@
    $x_{1}-x_{2}+x_{3}=0$
Assume that values $x_{1}=k_{1},x_{2}=k_{2},x_{3}=k_{3}$ are already chosen. Then the values for the rest of $x_{i}$ should be chosen, so that none of the following conditions are satisfied:

### v3

@@ Introduction
In [1], the matching arrangement of hyperplanes $MA(G)$ for a graph $G(V,E),|E|=n$ without loops and multiple edges was introduced as the arrangement in $\mathbb{R}^{n}$ that contains a hyperplane with the equation $x_{i_{1}}-x_{i_{2}}+...+(-1)^{k+1}x_{i_{k}}=0$ for every sequence of edges $(e_{i_{1}},...,e_{i_{k}})$ that form either a simple path of arbitrary length or a simple cycle of even length in $G$. The matching arrangement has the following property: if you arbitrarily select two vectors from the same region of the arrangement, use each vector in turn as a weight vector on the edges of graph $G$, and find the maximum-weight matching for each vector, then the maximum matching will be the same for both vectors [1]. In addition, there is a one-to-one correspondence between regions of the matching arrangement and LP-orientations of the matching polytope, a special polytope used in solving the maximum matching problem using linear programming methods [2].
In [3], Thomas Zaslavsky showed that the number of regions in a hyperplane arrangement depends on the partially ordered set of intersections of hyperplanes in the arrangement. Zaslavsky expressed this number through the characteristic polynomial of the hyperplane arrangement. Thus, if we calculate the characteristic polynomial of a hyperplane arrangement, we can find its number of regions.
@@
Let $A$ be any subspace arrangement in $\mathbb{R}^{n}$ defined over the integers and q be a large enough prime number. Then, $\chi_{A}(q)=|F_{q}^{n}\setminus\cup A|$. Equivalently, identifying $F_{q}^{n}$ with $\{0,1,...,q-1\}^{n}$, $\chi_{A}(q)$ is the number of points with integer coordinates in the cube $\{0,1,...,q-1\}^{n}$ which do not satisfy mod $q$ the defining equations of any of the subspaces in $A$.
The second section of this paper presents results on calculating the characteristic polynomials of the matching arrangement for some families of graphs using the finite field method. The definitions of a hyperplane arrangement, an intersection lattice and a characteristic polynomial that were used in the second section can be found in [6], which also contains the proof of the deletion-restriction lemma and the description of the finite field method that uses powers of prime.
@@
The fourth section of the paper provides a theoretical description of the Bolotnikov–Irmatov cryptosystem that uses the NP-completeness of the improper weight function problem.
The results from the second and the third sections are obtained by A.I. Bolotnikov. The cryptosystem in the fourth section is developed by A.I. Bolotnikov and A.A. Irmatov.
The characteristic polynomial of the matching arrangement for some classes of graphs
Definition 1.
Let $A$ be an arrangement of hyperplanes in $\mathbb{R}^{n}$. A partially ordered set (or poset) of intersections $L(A)$ is defined the following way: elements of $L(A)$ are the space $\mathbb{R}^{n}$ itself, and all nonempty intersections of hyperplanes in $A$, and $x\leq y$ in $L(A)$, if $x\supseteq y$.
Each element $x$ of the set $L(A)$ is an affine subspace of $\mathbb{R}^{n}$. Let $dim(x)$ be the dimension of $x$. If all hyperplanes of $A$ intersect in one point, then $L(A)$ is a lattice and is referred to as a lattice of intersections instead.
Definition 2.
Möbius function $\mu:P\times P\rightarrow\mathbb{Z}$ for a partially ordered set $P$ is defined by following conditions:
  1. 1.
    $\mu(x,x)=1$ $\forall x\in P$.
  2. 2.
    $\mu(x,y)=-\sum_{x\leq z<y}\mu(x,z)$ $\forall x<y\in P$.
  3. 3.
    $\mu(x,y)=0$, if $x$ and $y$ are not comparable.
If $P$ has a minimal element $\hat{0}$, then we write $\mu(x):=\mu(\hat{0},x)$.
Definition 3.
A characteristic polynomial of a hyperplane arrangement A is defined by the following formula
$\chi_{A}(t)=\sum_{x\in L(A)}\mu(x)t^{dim(x)}.$
@@
Let $G$ be a graph that consists of one cycle of length $n$, and let $\chi_{MA(G)}(x)$ be the characteristic polynomial of the matching arrangement $MA(G)$. Then,
@@
Let graph $G$ consist of one cycle of length $n$. We arbitrarily choose a vertex $A$ in graph $G$ and number the edges clockwise, counting the edges from $A$.
@@
Let us first consider the case when $n$ is even. Consider the linear operator $W:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ with the following matrix. The main diagonal of the matrix, as well as the next diagonal to the right, alternate between 1 and $-1$ (starting with 1), and there are 0s in all other places.
@@
The vector $t=(1,-1,...,(-1)^{k+1},0,...,0)$ is the normal vector to the hyperplane from the arrangement $MA(G)$, which corresponds to a path of length $k$ leaving $A$ and going clockwise. The operator $W$ maps this vector to vector $d_{k}=(0,...,0,1,0,...,0)$, in which 1 is at the $k$th position. For an arbitrary vertex $B$ of graph $G$, the vector $t_{B}=(1,-1,...,(-1)^{r+1},0,...,0)$, where $r$ is the length of the path from $A$ to $B$ clockwise, is normal to the hyperplane corresponding to the path from $A$ to $B$ in a clockwise direction. $t_{A}=(1,-1,...,1,-1)$ will denote the normal to the hyperplane corresponding to the entire cycle. Consider an arbitrary path in graph $G$ that leads clockwise from vertex $B$ to vertex $C$. If $B=A$, then the vector $t_{C}$ will be normal to the hyperplane corresponding to this path. If $C=A$ or the path does not pass through $A$, then $t_{C}-t_{B}$ will be normal to the corresponding hyperplane. If the path passes through $A$ but does not start or end there, then $t_{A}-t_{B}+t_{C}$ will be normal to the corresponding hyperplane. Operator $W$ maps $t_{B}$ to $d_{r}$, $t_{C}-t_{B}$ to $d_{k_{2}}-d_{k_{1}}$, where $k_{2}>k_{1}$, and $t_{A}-t_{B}+t_{C}$ to $d_{n}-d_{k_{1}}+d_{k_{2}}$, where $k_{1}>k_{2}$.
@@
    $\forall i_{1},i_{2},1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}+x_{i_{1}}=0$.
Note that since there is the invertible linear operator $W$ that maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of $S$, their intersection lattices are isomorphic and their characteristic polynomials are equal. We will compute the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number $p$, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
@@
    $\forall i_{1},i_{2},1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}+x_{i_{1}}=0$.
We will find the number $\chi_{S}(p)$ combinatorially. First, we choose a value for $x_{n}$. This can be done in $p-1$ ways. Let us say we chose $x_{n}=k$, $k\neq 0$. Then, the remaining $x_{i}$ should be selected so that they do not satisfy any of the following conditions:
@@
Let us choose a weak ordered partition $\pi=(B_{1},...,B_{p-n})$ of the set $\{1,...,n-1\}$ into $p-n$ blocks (“weak” means that some of the blocks can be empty), that is $\bigcup B_{i}=\{1,...,n-1\}$ and $B_{i}\cap B_{j}=\emptyset,i\neq j$. Arrange all the elements of $F_{p}$ in a row as follows: $0,k,2k,...,(p-1)k$. The first two elements of this row are “locked”: none of the $x_{i}$ can take these two values. Starting with the third element in this row, elements of $B_{1}$ are placed one by one in the descending order onto the elements of the row without any gaps. Once all elements of $B_{1}$ are placed, a gap with the size of one element of the row is made, and elements of $B_{2}$ are placed on the following elements of the row in the same descending order. We keep placing elements of $B_{i}$ onto the elements of the row in the descending order and making one-element gaps when moving from $B_{i}$ to $B_{i+1}$, until all elements from $\{1,...,n-1\}$ are placed.
Now for every $i$, $10\leq i\leq n-$ 1, we choose the element of the row, in which we placed $i$, as $x_{i}$. Conditions $x_{i}=0$ and $x_{i}=k$ aren’t satisfied for any $i$, because the first two elements of the row were left “locked”. Condition $x_{i_{2}}-x_{i_{1}}=0$ is not satisfied for any $i_{1}$ and $i_{2}$, because we put no more, then one element of $\{1,...,n-1\}$ onto each element of the row. Condition $0\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}-x_{i_{1}}=k$ is not satisfied for any $i_{1}\leq i_{2}$, because for any $B_{i}$ the elements of $B_{i}$ were placed in the descending order, and there was a one-element gap between elements of $B_{i}$ and $B_{i+1}$. Therefore, a vector $(x_{1},...,x_{n-1})$ is constructed for a given partition $\pi$, and it does not satisfy any of the four conditions.
Conversely, given a vector $(x_{1},...,x_{n-1})$ that does not satisfy any of the four conditions, a placement of $\{1,...,n-1\}$ onto the row $0,k,2k,...,(p-1)k$ can be recovered, and in turn a partition $\pi$ can be recovered from this placement. This means that the number of vectors that do not satisfy any of the four conditions is equal to the number of weak ordered partitions of $\{1,...,n-1\}$ into $p-n$ blocks. This number is equal to the number of functions $f:[n-1]\rightarrow[p-n]$. There are exactly $(p-n)^{n-1}$ of such functions. And considering the choice of $x_{n}$ on the first step, the resulting formula for the characteristic polynomial of $S$ is $\chi_{S}(p)=(p-1)(p-n)^{n-1}$. Thus, in case when graph $G$ consists of one cycle of even length, $\chi_{MA(G)}(x)=(x-1)(x-n)^{n-1}$.
Now let $n$ be odd. Let $L$ be the following arrangement of hyperplanes: $L=MA(G)\cup\{H\}$, where $H$ is a hyperplane with the equation $x_{1}-x_{2}+...-x_{n-1}+x_{n}=0$.
Let the linear operator $W:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ be defined similar to the operator in the case of an even cycle. The main diagonal of the matrix, as well as the next diagonal to the right, alternate between 1 and $-1$ (starting with 1), and there are 0s in all other places. We will use $d_{k}$ and $t_{B}$ to denote the same vectors as in the case of an even graph, with $t_{A}=(1,-1,...,1,-1,1)$ now being a normal vector to $H$. The representations of normal vectors to hyperplanes of $L$ as linear combinations of vectors $t_{A}$, $t_{B}$, and $t_{C}$ are mostly the same, the only difference is that if the path from $B$ to $C$ passes through $A$ but does not start or end there, then $t_{A}-t_{B}-t_{C}$ will be normal to the corresponding hyperplane.
@@
    $\forall i,1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}-x_{i_{1}}=0$.
Note that since there is the invertible linear operator $W$ that maps the set of normal vectors to hyperplanes of $L$ to the set of normal vectors to hyperplanes of $S$, their intersection lattices are isomorphic and their characteristic polynomials are equal. We will compute the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number $p$, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations:
@@
    $\forall i,1\leq i_{1}\leq i_{2}<n:x_{n}-x_{i_{2}}-x_{i_{1}}=0$.
We will find the number $\chi_{S}(p)$ combinatorially. First, we choose a value for $x_{n}$. This can be done in $p-1$ ways. Let us say we chose $x_{n}=k$, $k\neq 0$. Then, the remaining $x_{i}$ should be selected so that they do not satisfy any of the following conditions:
@@
    $\forall i_{1},i_{2},1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}+x_{i_{1}}=k$.
Let us consider two possible cases.
Case 1: $\forall i,0\leq i\leq n-1,\leavevmode\nobreak\ x_{i}\neq k/2$. In this case, we can consecutively choose values for all $x_{i}$. There are $p-3$ possible values for $x_{1}$. After a value is chosen for $x_{1}$, there are $k-5$ possible values for $x_{2}$: if value $t$ is chosen for $x_{1}$, then we cannot choose $t$ and $k-t$ for $x_{2}$. After a value is chosen for $x_{2}$, there are $p-7$ possible values for $x_{3}$, and so on. In total, in this case there are $(p-3)(p-5)(p-7)...(p-(2n-1))$ possible ways to choose values for $x_{1},...,x_{n-1}$.
Case 2: for some $i$ $x_{i}=k/2$. In this case, we start with choosing in $n-1$ ways, which of $x_{i}$ is equal to $k/2$, and then for the rest of $x_{i}$ the values are chosen the same way, as in Case 1. Therefore, in this case there are $(n-1)(p-3)(p-5)...(p-(2n-3))$ ways to choose values for $x_{1},...,x_{n-1}$.
@@
Let $\hat{S}$ be the arrangement $\hat{S}=S\setminus{\hat{H}}$, where $\hat{H}$ is the arrangement with the equation $x_{n}=0$. The invertible linear operator $W$ maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of $\hat{S}$ , which means that their intersection lattices are isomorphic and their characteristic polynomials are equal.
The characteristic polynomial of $\hat{S}$ can be found with the use of the deletion-restriction lemma: $\chi_{\hat{S}}(x)=\chi_{S}(x)+\chi_{S_{\hat{H}}}(x)$, where $S_{\hat{H}}$ is the arrangement of hyperplanes in $\hat{H}$ that consists of intersections of $\hat{H}$ with other hyperplanes of $S$. The following are the equations of hyperplanes of $S_{\hat{H}}$:
@@
    $\forall i_{1},i_{2},1\leq i_{1}\leq i_{2}\leq n-1:x_{i_{2}}+x_{i_{1}}=0$.
We will find $\chi_{S_{\hat{H}}}(x)$ using the finite field method. For a sufficiently large prime number $p$, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations above. Let us choose values for $x_{i}$ consecutively. There are $p-1$ ways to choose a value for $x_{1}$, there are $p-3$ ways to choose a value for $x_{2}$ after choosing a value for $x_{1}$, and so on. Therefore, $\chi_{S_{\hat{H}}}(p)=(p-1)(p-3)(p-5)...(p-(2n-3))$.
After applying the deletion-restriction lemma, the following result is obtained: in case when $G$ consists of one cycle of odd length, $\chi_{MA(G)(x)}=(x-1)(x-3)(x-5)...(x-(2n-3))(x-(n-1))$. ∎
@@
Let graph $G$ consist of two subgraphs that intersect in one vertex – one subgraph is a complete graph $K_{3}$ and the other is the “tail” – a path of an arbitrary length. Then, $\chi_{MA(G)}(x)=(x-1)(x-3)(x-4)...(x-(n-2))(x-(n-1))(x-(n-1))(x-n)$.
@@
Let graph $G$ consist of two subgraphs that intersect in one vertex – one subgraph is a complete graph $K_{3}$ and the other is the “tail”– a path of an arbitrary length. Let us number the edges of $G$ in the following order: first we number the edges of $K_{3}$, then we number the edges of the tail, starting with the edge that is adjacent to $K_{3}$.
@@
Let $W$ be a linear operator $W:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ with the matrix shown on Fig. 3.
@@
Let $d_{1}=(1,0,...,0)$, $d_{2}=(0,1,0,...,0)$, $d_{3}=(0,1,-1,0,...,0)$, $d_{i}=(0,0,0,1,-1,...,(-1)^{i},0,...,0)$, where $(-1)^{i}$ stands on the $i$th place, $i>3$. Operator $W$ maps $d_{i}$ into a vector $(0,...,0,1,0,...,0)$, where 1 stands on the $i$th place.
Let us consider three possible cases for paths in $G$.
@@
Case 3: The path goes from $K_{3}$ into the tail. The following are all possible cases of the intersection of this path and the $K_{3}$ subgraph. If this intersection is $\{e_{1}\}$, then $d_{i}-d_{1}$ is a normal vector to the corresponding hyperplane. If $\{e_{2}\}$ is the intersection, then $d_{i}-d_{2}$ will be the normal vector. If $\{e_{2},e_{3}\}$ is the intersection, then $d_{i}-d_{3}$ will be the normal vector. If $\{e_{3},e_{1}\}$ is the intersection, then $d_{i}-d_{1}+d_{2}-d_{3}$ is the normal vector.
Let $S$ be the arrangement that consists of the following hyperplanes:
@@
    $x_{1}-x_{2}+x_{3}=0$.
Note that since there is the invertible linear operator $W$ that maps the set of normal vectors to hyperplanes of $MA(G)$ to the set of normal vectors to hyperplanes of $S$, their intersection lattices are isomorphic and their characteristic polynomials are equal. We will compute the characteristic polynomial $\chi_{S}(x)$ using the finite field method. For a sufficiently large prime number $p$, $\chi_{S}(p)$ is equal to the number of vectors $(x_{1},...,x_{n}),x_{i}\in F_{p}$, that do not satisfy any of the equations
@@
    $x_{1}-x_{2}+x_{3}=0$.
Assume that values $x_{1}=k_{1},x_{2}=k_{2},x_{3}=k_{3}$ are already chosen. Then, the values for the rest of $x_{i}$ should be chosen, so, that none of the following conditions are satisfied